### v1

# Large Segmentation Design of Robust Photonic Quantum Gates

Abstract
We analyze the fidelity and power loss of photonic quantum gates, implemented by a segmentation of coupled optical waveguides, as functions of the number of segments. We show that designs with a large number of segments mitigate effectively the systematic errors for a wide range of error correlation strength and variance. We quantify the design roughness effect on the power loss and analyze the performance of various methods for smoothing the configuration of the waveguides. We observe an asymptotic scaling behaviour of the gate fidelity and power loss with the number of segments.
I Introduction
The quest for harnessing the potential of quantum phenomena has been a driving force behind intensive fundamental research and technology efforts, including Quantum Information Processing (QIP) bennett2000quantum .
Quantum computers hold the promise for an exponential speedup in solving certain complex computational problems and a capability to decrypt classical encryption protocols in real-time, thereby opening new frontiers in computing and cryptography. Despite substantial theoretical advancements in quantum algorithms QuantumAlgorithms1 ; QuantumAlgorithmSurvey , the practical realization of a quantum information processor remains a formidable challenge.
One of the paramount obstacles in achieving practical quantum computation lies in the noise and errors that affect the performance of the fundamental building blocks of quantum algorithms: the qubit’s state preparation, the quantum gates, and the measurement process Nielsen2000 . These errors and noise have various sources, such as fabrication imperfections Kyoseva2019 , inaccurate coupling strengths and the interaction with the classical environment noiseoneoverf . In response to these challenges, two primary routes have been taken: error-correcting codes QEC1 and error mitigation schemes NoiseMitigation2 ; QNoiseMitigation1 .
Error-correcting codes construct logical qubits using additional entangled physical qubits and can significantly outperform their physical counterparts. Notable among these codes are stabilizer codes, exemplified by the surface code Fowler2012 , which exhibits a relative tolerance to local errors. Fault-tolerant quantum computation using error correcting codes requires the fidelity of the physical quantum gates to be above a certain threshold, making the reduction of the physical errors to be of utmost importance NoiseThreholdQC ; QuantumErrorCorrection .
Error mitigation schemes seek to reduce errors using sophisticated implementation and techniques, for a review see QNoiseMitigation1 ; NoiseMitigation2 . One such framework is the segmented composite pulse design, which was originally used in atomic physics and nuclear magnetic resonance, employing a combination of constant pulses to minimize errors during the system evolution CPNMR ; CPNMR2 ; CPAtomicPhysics ; CPAtomicPhysics2 . This error mitigation technique holds a promise for improving the performance of complex quantum algorithms, and has a broad applicability in diverse quantum computing realizations of quantum information processing devices.
In SmoothCP smooth designs were proposed in a general composite pulses scheme, and HUSAIN2013145 suggested symmetric and antisymmetric composite pulses with a large number of pulses in NMR. It has been recently shown in Kaplan2023 that this method can be implemented in Silicon photonics KLM for designing robust high fidelity quantum gates by varying the waveguide geometry, and specifically its cross-section. The focus in Kaplan2023 has been on having a small number of segments, under the assumption of fully correlated noise of the different waveguides. It was proven in YaronIdoNoiseThreshold that there is a correlation strength threshold for the segmented design to be an effective error mitigating scheme.
The aim of the current work is to analyze the composite pulse scheme for optimizing robust designs with a large number of segments. We consider both the fidelity of the quantum gates as well as the power loss of the designs for varying systematic noise correlations. Applied to the photonic dual-rail realization, our methods yield robust high-fidelity unitary designs for single-qubit and hence also for the probablistic double-qubit gates such as CZ and CNOT Knill2001 . We consider different manufacturing configurations in Silicon and Lithium Niobate, as Silicon is widely used in this field, while the Lithium Niobate is considered to be a promising material for designing active QIPs.
Our main contributions are :
  1. 1.
    Enhanced Quantum Gate Robustness: We demonstrate numerically that increasing the number of composite segments $N$ results in a substantial improvement in the quantum gate mean fidelity $\mathbf{E}\left[F\right]$, and a reduction in its standard deviation $\sigma_{F}\left[F\right]$ with a certain scaling with $N$. We further show that $N\sim O(10)$ is effective for most practical cases.
  2. 2.
    Power Loss Analysis: We quantify the impact of increasing $N$ on the power loss in the quantum gate due to the back-scattering effects in the segments. We argue analytically and show numerically that the power loss grows linearly with the number of segments, which makes it necessary to apply smoothing methods to the large $N$ designs.
  3. 3.
    Correlation Strength: We study the impact of error correlation strength between the different segments on the gate fidelity. We show that enhancing the correlation strength among the errors in different segments leads to a significant increase in the expected fidelity of the quantum gate, hence providing valuable insights to strategies for choosing the optimal number of segments given a correlation strength. For a given required gate robustness, larger correlation strength implies that less segments are needed.
  4. 4.
    Adiabatic and Fresnel Designs: We introduce smooth segmented gates designs based on the principles of the Adiabatic and Fresnel scattering, that exhibit a reduced power loss while keeping the same fidelity performance. This offers a complementary tool to the traditional segmentation method.
The paper is organized as follows: In Section II, we briefly review the composite pulse design of quantum gates and the error model in the optical waveguides framework. In Section III we present the ingredients of the cost function optimization framework including the optimization range, the quantum gate fidelity and the smooth design conditions for minimizing the power loss. In Section IV we show the optimization results for the gates fidelity and power loss for a varying number of segments and for different smooth designs. Section V is devoted to a discussion and conclusions. Additional details are provided in four appendices.
II Segmented Composite Pulses
Figure 1: We divide each of the two waveguides to $N$ segments. Each segment $i$ is characterized by its coupling $\kappa_{i}$, mode-mismatch coefficient $\Delta_{i}$ and length $z_{i}$. The interpolation map for $\kappa$ and $\Delta$ as a function of the waveguides widths $W_{1},W_{2}$, in $nm$ units. These results are extracted from a Lumerical FDTD Eigensolver simulation, for Silicon on insulator waveguides platform with a height of $220nm$ and a gap of $1\mu m$. One of the waveguides was illuminated with $1.55\mu m$ light source.
II.1 Robust Composite Quantum Gates
The presence of noise within the physical two-level system leads to deviations from the desired quantum gate operation. To quantify the extent to which the noisy gate deviates from the ideal gate, we define the gate fidelity $F$ by the trace norm:
$F\left(U_{\text{ideal}},U(\boldsymbol{\epsilon})\right)=\frac{1}{2}\left|\text{Tr}\left(U_{\text{ideal}}^{\dagger}U(\boldsymbol{\epsilon})\right)\right|\ ,$ (1)
where $U_{\text{ideal}}$ represents the desired ideal unitary gate, and $U(\boldsymbol{\epsilon})$ represents its physical realization, which depends on a set of $m$ jointly distributed random errors denoted as $\boldsymbol{\epsilon}=\{\epsilon^{a}\}_{a=1}^{m}$. The fidelity is a random function that takes values in the interval range $[0,1]$, where zero and one correspond to the absence of errors and to the maximal deviation from the desired quantum gate operation, respectively. The primary objective of our analysis is to maximize the mean fidelity, denoted as $\mathbf{E}\left[F\right]$, and to minimize its standard deviation, denoted as $\sigma[F]$:
$\mathbf{E}\left[F\right]$ $=$ $\mathbb{E}_{\boldsymbol{\epsilon}}\left[F\left(U_{\text{ideal}},U(\boldsymbol{\epsilon})\right)\right]\ ,$
$\sigma[F]$ $=$ $\sqrt{\mathbb{E}_{\boldsymbol{\epsilon}}\left[F\left(U_{\text{ideal}},U(\boldsymbol{\epsilon})\right)^{2}\right]-\mathbf{E}\left[F\right]^{2}}\ .$ (2)
The composite pulses is a framework for error mitigation where $U(\boldsymbol{\epsilon})$ is realized as a product of $N$ physical unitary gates:
$U_{cp}(\vec{\boldsymbol{\epsilon}})=\prod_{i=1}^{N}U_{i}(\boldsymbol{\epsilon}_{i})\ ,$ (3)
where $\vec{\boldsymbol{\epsilon}}=(\boldsymbol{\epsilon}_{1},...,\boldsymbol{\epsilon}_{N})$, $\boldsymbol{\epsilon}_{i}=\{\epsilon^{a}_{i}\}_{a=1}^{m}$, and $U_{i}$ are constructed to achieve higher fidelity. We will consider this error mitigation scheme in the framework of dual-rail waveguides in integrated photonics.
II.2 Coupled Optical Waveguides
Consider the following physical photonic realization of single-qubit quantum gates, which is also the basis for constructing probabilistic two-qubit gates AmnonYariv . A coupled system of two waveguides can be described by two parameters: the coupling coefficient $\kappa$ and the mode-mismatch coefficient $\Delta$ (known also as the detuning coefficient). The coupling coefficient is given by the integral over the spatial modes-overlap:
[…]
where $(1)$ and $(2)$ refer to the two waveguides in the dual-rail representation, $(x,y)$ are the transverse directions, $z$ denotes the propagation direction and $\epsilon(x,y)$ and $\epsilon^{(2)}(x,y)$ are the permittivity distributions of the coupled waveguides system and of waveguide $(2)$ alone, respectively. The mode-mismatch coefficient is determined by the difference between the propagation constants of the two waveguides:
$\Delta=\frac{\beta_{1}-\beta_{2}}{2}\equiv\frac{2\pi}{\lambda}\left(n_{{eff}_{1}}-n_{{eff}_{2}}\right)\ ,$ (5)
where $n_{{eff}_{i}}$ are the effective refraction indices in the first and second waveguides, respectively, and depend on the geometry of the waveguides. A simple solution for the propagation differential equation is constant coupling and mode-mismatch, which we refer to as one segment. It is given by (see appendix A):
$U\left(\kappa,\Delta,z\right)=\cos\left(\Omega_{g}z\right)\mathbf{I}-i\frac{\kappa}{\Omega_{g}}\sin\left(\Omega_{g}z\right)\mathbf{X}$
$+i\frac{\Delta}{\Omega_{g}}\sin\left(\Omega_{g}z\right)\mathbf{Z}\ ,$ (6)
where $\Omega_{g}\equiv\sqrt{\Delta^{2}+\kappa^{2}}$, and $\mathbf{I},\mathbf{X},\mathbf{Z}$ are the identity matrix and Pauli matrices, respectively.
II.3 Segmentation
We will analyze quantum gates designs that consist of $N$ segments (see Fig. 1). Supported by experimental evidence ExperimentCP , we consider the case where in each segment, the noise affects both waveguides identically (full correlation), i.e. $\epsilon^{(1)}_{k}=\epsilon^{(2)}_{k},~{}k=1,...,N$. This implies a possible significant error in the coupling coefficients $\kappa_{k}$, but affects less the mode-mismatch coefficients $\Delta_{k}$. As to the correlation between the errors of the different segments, we will initially assume a full correlation between them, and will later relax this condition and analyze the case of a general correlation matrix. The $N$-segmented gate reads:
$U^{(N)}=\prod_{k=1}^{N}U(\kappa_{k},\Delta_{k},z_{k},\epsilon_{k})\ ,$ (7)
where $z_{k},\kappa_{k},\Delta_{k},\epsilon_{k}$ are the length of the $k^{th}$-segment, its coupling, its mode-mismatch coefficient and its noise, respectively.
II.4 Power Loss
Implementation of the composite pulse error mitigation scheme in Silicon photonics or in any other physical platform is accompanied by power loss due to back-scattering. This back-scattering is a direct consequence of the non-smooth profile of the segmented waveguide’s width, and captures the total power loss (see Appendix B).The gate reflection power loss is total power transmission coefficient $P_{t}$:
$P_{t}\equiv T=\prod_{k=1}^{N-1}T_{k}(\kappa_{k},\Delta_{k},z_{k},\epsilon_{k})\ ,$ (8)
where $T_{k}$ is the power transmission coefficient from the segment $k$ to the segment $k+1$.
Specifically, in photonics, we can estimate the back-scattering by the Fresnel equation that governs the reflection and transmission coefficients of a light transmitted between two different materials with refractive indices $n_{1}$ and $n_{2}$ (we consider the intensity ):
$R=\left(\frac{n_{1}cos\theta_{i}-n_{2}cos\theta_{t}}{n_{1}cos\theta_{i}+n_{2}cos\theta_{t}}\right)^{2}\ ,$ (9)
where we chose TE polarization and $\theta_{i}$ and $\theta_{t}$ are the incident and transmitted light angle, respectively. In guided propagation, the refractive index is replaced by $n_{eff}$, which is calculated from the overlap integral of the spatial cross-section field profile with the cross-section refractive index.
For our analysis, focusing on the TE polarization (the basic mode) and assuming small angles related to the $z$ axis, we evaluate the total total reflection coefficient as:
$R=1-T=1-\prod_{i=1}^{N-1}\left(1-\left(\frac{n_{eff_{i+1}}-n_{eff_{i}}}{n_{eff_{i+1}}+n_{eff_{i}}}\right)^{2}\right)\ .$ (10)
III Optimization Framework
Our optimization search method is subject to several constraints. While our goal is to obtain high fidelity quantum gates, we aim to achieve it with as smooth as possible design in order to minimize the power loss. As discussed, abrupt changes in the waveguide’s width increase the power loss due to back scattering, and also give rise to higher spatial modes in the waveguide that distort the signal, hence changing the qubit state. Moreover, waveguide designs must be physically implementable subject to the manufacturing resolution.
We divide each waveguide into a sequence of $N$ small segments, each of length $z_{i}$ as shown in Fig. 1. The two waveguides are separated by a gap (defined as the distance between the centers of the waveguides) and have a height $h$. Our optimization uses the Lumerical FDTD results as a look-up table to inverse-map the geometry to the coupling coefficient and the mode-mismatch parameters.
III.1 Fidelity
We will assume that the errors in the two waveguides widths $W_{1}$ and $W_{2}$ are fully correlated and can be sampled from a Normal distribution with zero mean and a standard deviation $\sigma$, $D\sim\mathcal{N}(0,\sigma)$. For our analysis, we consider the dimensionless parameter $\sigma_{\%}\equiv 100\frac{\sigma}{W_{mean}}$, where
$W_{mean}\equiv\frac{1}{2N}\sum_{i}\left(W_{i}^{(1)}+W_{i}^{(2)}\right)\ ,$ (11)
and we focus on the range $1\%-10\%$, which is approximately around $10_{nm}-50_{nm}$. The fidelity loss term reads:
$\mathcal{L}_{F}=1-E_{\delta W\sim D}\left[\frac{1}{2}Tr\left(U_{ideal}^{\dagger}U\right)\right]\ ,$ (12)
where $U_{ideal}$ is the ideal quantum gate and $U$ is its physical realization.
III.2 The Range of Width
We add a regularization in order to restrict the width of the waveguides to be in a range between the minimum and the maximum of an allowed region:
[…]
$+$ $\min{\left(W_{min}-W_{i},0\right)^{2}})\ \bigg{)}\ ,$
where $\lambda_{W}$ is a hyperparameter for the minimal and maximal widths. We define the optimization domain $\Omega=(\Omega_{dz},\Omega_{w})$ as the set of all experimentally realizable waveguide widths. The number $N$ of segments in the design varies between 1 and 200, and we denote by $L$ the total length of the design. The degrees of freedom for the optimization are the set of widths and lengths of each segment: $\left\{W^{\text{L}}_{i},W^{\text{R}}_{i},dz_{i}\right\}_{1\leq i\leq N}$.
The waveguide width initialization is sampled uniformly in the range between the maximum and minimum of the feasible domain. Additionally, strong regularization is activated once the design is out of the optimization domain $\Omega$. It is also possible to add another regularization for the total design length $L$, but in our analysis we do not impose this restriction. Almost all the design lengths are between $10\mu m$ and $100\mu m$.
III.3 Smooth Design Conditions
III.3.1 Baseline optimization
We define the baseline optimization as a search for the highest fidelity designs even with having abrupt changes between segments cross-section widths. This exhibits scattering and reflections, higher spatial modes generations and power loss, and requires a very high precision for manufacturing. Designs will be compared to the baseline ones.
III.3.2 Minimizing Fresnel Reflection
In order to reduce the power loss, we add a regularization term which reduces the back reflection coefficient. Thus, among the robust high-fidelity design parameter space, we search for the designs with the minimal power loss. The total cost function reads:
$\mathcal{L}=\mathcal{L}_{F}+L_{\Omega}+\lambda_{R}R\ ,$ (14)
where $\mathcal{L}_{F},L_{\Omega},R$ are given by (12), (13) and (10), respectively.
III.3.3 Imposing An Adiabaticity Condition
The adiabatic theorem guarantees, under certain conditions, a smooth transition between the ground states of two different Hamiltonians adiabatic . In our setup we impose an adiabatic condition on the design in the form:
$\left|\frac{\partial\beta}{\partial z}\right|\ll\beta(z)\ .$ (15)
Assuming a power law dependence of $\beta$ on the cross-section area: $\beta=aW^{m}$ for some $a$ and $m$, where $W$ is the width of the waveguide, we get:
$\left|\frac{\partial W/\partial z}{W(z)}\right|\ll 1\ .$ (16)
We define a corresponding regularization term:
$A=\max_{i}\frac{\left|W_{i+1}-W_{i}\right|}{\left(\frac{W_{i+1}+W_{i}}{2}\right)\left(\frac{z_{i+1}+z_{i}}{2}\right)}\ .$ (17)
The cost function now reads:
$\mathcal{L}=\mathcal{L}_{F}+L_{\Omega}+\lambda_{A}A\ .$ (18)
Figure 2: Example of two 100-segments designs of $\sqrt{X}$ gate in Silicon, with width noise standard deviation of $\sigma=30$nm ($\sigma_{\%}\approx 5\%$). Top: the Fresnel-based design with expectation fidelity of $0.91$ and reflection coefficient of ±$\pm 10^{-3}$. Bottom: the baseline with expectation fidelity of $0.9$ and reflection coefficient of ±$\pm 10^{-2}$. The roughness standard deviation (width change along the design) is $23$nm and $76$nm, respectively.
III.4 Error Correlations
The errors at different segments may be partially correlated. In order to quantify the effect on the design performance of different strength of correlations between the errors of different segments we introduce the error correlation matrix between the $k,l\in\{1,...N\}$ segments:
$\rho_{kl}(\mu)=\begin{cases}1&k=l\\ \frac{1}{\left(\left|k-l\right|+1/2\right)^{2\mu}+1}&k\neq l\ .\end{cases}$ (19)
The dimensionless parameter $\mu$ controls the strength of the correlation: $\mu=0$ implies strong correlations between segments while $\mu\rightarrow\infty$ implies no correlation. The results of our correlations simulation will be discussed in the the simulation section (see IV.3).
III.5 Waveguide Material
We optimize the design for two types of materials: Silicon, which is widely used in photonic applications and has a relatively high refractive index (about 3.48), and LiNbO3 which is considered to be a promising candidate for photonic applications due to its outstanding optical properties, including a wide transparency window $(0.35\mu m-5\mu m)$, large optical nonlinearity and electro-optic coefficient, and strong acousto-optic effect. The Lithium Niobate has negative uniaxial birefringence, but we assume that only the first $TE_{01}$ mode is occupied.
For both materials we use a gap of $1.0~{}\mu m$ and the width of each waveguide is varied between $500~{}nm$ to $850~{}nm$. We further assume manufacture resolution (smallest resolution in the optimization) to be $20~{}nm$, waveguide height of $220~{}nm$, etch of $150~{}nm$ and wavelength of $1.55~{}\mu m$. To evaluate the mode-mismatch and coupling coefficients of the design we use Lumerical Eigensolver simulation (as shown in Fig. 1). We simulate different design geometries and build a look-up table to express the mode-mismatch coefficient $\Delta$ and the coupling coefficient $\kappa$ in terms of the geometrical waveguide parameters. The resolution of the geometry is $5~{}nm$, and we use a bi-linear interpolation to evaluate the values in-between.
IV Optimization Results
In the optimization we divide each waveguide into $N=100$ segments. We use the PyTorch Package, with Adam optimizer, CosineAnnealing LR scheduler, with an initial learning rate of 0.1, batch size of 1024 samples, and the number of epochs was set to be around 5000. The device length is initialized randomly between $25~{}\mu m$ to $100~{}\mu m$. In our analysis we focus on the Pauli $X$ (without mode-mismatch) and the Hadamard $H$ (with mode-mismatch) single-qubit quantum gates. We found that other single-qubit gates show a similar behaviour. In Fig. 2 we see examples of composite designs (baseline and Fresnel based). Diagram of the trajectories on the Bloch sphere of designs for different optimization methods are shown in Fig. 3.
Figure 3: Bloch Sphere diagram of two optimization results. Left - Hadamard gate with 10 segments, initial state is $|0\rangle$ and final state is |+ket|$\rangle$| + ⟩. Right - Pauli X gate with 50 segments, initial state is $|+\rangle_{y}$ and final state is $|-\rangle_{y}$.
IV.1 Mean Fidelity and Variance
Fig. 4 shows the expectation value and standard deviation of the fidelity for different numbers of segments taken in the range 1 to 200. To clean the optimization numerical noise, we avergated over 20 optimizations (for each number of segments). As can be clearly seen from the plot, increasing the number of segments leads to an oncrease in the gate fidelity, and the most significant improvement is obtained when the increasing the number of segments from 1 to 10. In Fig. 5 we plot the expectation value of the fidelity for different error magnitudes. Empirically, we find that the fidelity for Adiabatic and Fresnel designs are very similar to the baseline (see Fig. 9 and Appendix C for the list of the designs). We numerically observe large $N$ scaling laws for the expectation value and variance of the fidelity as functions of the number of segments:
$\mathbf{E}\left[F\right]\sim 1-\frac{a}{N},~{}~{}~{}~{}\sigma[F]\sim N^{-b}\ ,$ (20)
for some constants $a$ and $b$, which we expect to be non-universal and depend on the geometry of the device, the material and the wavelength. In our example, they are $a=0.1$, and $b=0.38$.
Figure 4: Mean Fidelity and standard deviation (shaded area) versus the number of segments for Pauli X, gate and Litium-Niobit material, assuming Normally distributed width noise with $\sigma=30$nm ($\sigma_{\%}\approx 5\%$). We see that the mean fidelity increases and variance decreases as the number of segments grows. This plot was created using an ensemble average over 20 optimizations for each $N$.
Figure 5: The expectation value of the fidelity for different number of segments, for increasing noise width. The design is optimized for $\sigma_{\%}\approx 5\%$. Increasing the number of segments leads to a higher robustness for large width errors.
IV.2 Power Loss
To evaluate the power loss of the design, we use Lumerical simulator to measure the transmitted light for each transition between two neighbouring segments with different widths. In appendix B we present a comparison between different power loss calculation methods (Fresnel coefficients, FDTD, varFDTD and EME simulations), while here we focus on the comparison between the Fresnel coefficient (up to a global factor, see appendix B for details), and the FDTD simulation result (which is considered the most accurate among the simulation methods).
Fig. 6 shows the power loss for different number of segments, for the Hadamard gate made of Litium-Niobit material. While there are some differences between the Fresnel coefficients and the FDTD, their overall behaviour is very similar. We find empirically that optimizing over the Fresnel coefficients leads to a reduction in the total power loss estimated by the FDTD. To calculate the total power loss per design, we converted the loss per transition (between segments) using Eq. 8. Fig. 6 shows the averaged result over 10 optimizations.
Figure 6: Power loss versus number of segments for Hadamard, Litium-Niobit, using Fresnel coefficients (up to a global scale factor) in dashed line and FDTD simulation in solid line. We see an increase in the power loss as the number of segments increases, with an improvement when employing the Fresnel and the Adiabatic smoothing methods. This plot is created using an ensemble average over 10 optimizations for each $N$ and optimized for $\sigma_{\%}\approx 5\%$.
The averaged total power loss scaling is asymptotically linear in $N$:
$\lim_{N\rightarrow\infty}\langle P_{t}\rangle=\lim_{N\rightarrow\infty}\left(1-\left(1-\langle R_{avg}\rangle\right)^{N-1}\right)=NR_{avg}\ ,$ (21)
where $R_{avg}$ is the average reflection coefficient of a transition between two adjacent segments. We plot in Fig. 7 the power density transmission cross-section curve using FDTD Lumerical simulator.
Figure 7: Example of power density transmission cross-section curve using FDTD Lumerical simulator. The designs are of Silicon waveguides with $N=3$: Top - Baseline, Bottom - Fresnel-based. Green - Injected light in the beginning of the waveguide, blue - transmitted light to the end of the waveguide. The horizontal axis is the spatial cross-section of the waveguide, the vertical axis is the electromagnetic wave power density profile.
IV.3 Fidelity and Error Correlation
When the correlation between the errors at different segments decreases, the composite pulse scheme for error mitigation becomes less effective YaronIdoNoiseThreshold . Here we analyze the impact of having a large number of segments when the error correlation decreases, where we model the error covariance matrix as (19). We plot in Fig. 8 the mean value and the standard deviation of the fidelity versus the error correlation strength $\mu$ (left, full correlation, and right, no correlation), for different numbers of segments for Pauli X gate, Silicon material. As expected, designs with a large number of segments show a better performance when the error correlations are small. We observe the same behaviour for other quantum gates.
Figure 8: Mean value and standard deviation of the fidelity versus the error correlation strength (left - full correlation and right - low correlation), for different number of segments, for gate X, Silicon. We see the improvement of the design performance when increasing the correlation and the number of segments. This plot was created using an ensemble average over 10 optimizations for each $N$ and optimized for $\sigma_{\%}\approx 5\%$.
IV.4 Fidelity Versus Power Loss - An Effective Error Mitigation Region
Taking into account both the fidelity of the quantum gate and the power loss of the design, we plot in Fig. 9 the mean fidelity versus the power loss and identify the effective error mitigation region. The designs are listed in table 1 in appendix C
Figure 9: Fidelity versus power loss (As calculated with the FDTD simulation). Each point represents a single design, the color indicates the number of segments $N$, the shape indicates the optimization type, and the size of the point indicates the fidelity’s standard deviation. The dashed line defines a domain that includes designs, where the mean fidelity is very high and the power loss is very low. Table 1 in appendix C gives the details of the designs.
V Discussion and Outlook
We studied the performance of the segmented composite pulse error mitigation scheme applied to the design of photonic quantum gates, when increasing the number of segments $N$ and varying the error correlation strength $\mu$. We considered the fidelity and the power loss of single-qubit quantum gates realized by optical waveguides in the dual rail framework. We quantified the impact of increasing $N$ on the fidelity and the power loss of the quantum gates and observed certain asymptotic scaling laws. Empirically we found that $O(10)$ number of segments yields effective designs.
To reduce the power loss when increasing $N$, we applied two optimization regularizations by imposing constraints on the Fresnel reflection coefficient and an adiabatic condition on the waveguides width profile. We have not identified a significant difference in the gate performance between the two regularizations. However, from practical considerations it is simpler to apply the adiabatic one since it does not require a pre-calculation of the effective refraction index in the waveguide.
We studied the effect of varying the correlation strength between the errors of different segments. Large correlations between the errors imply better gate fidelity, since the different segments can effectively compensate the accumulated errors. We showed that a large number of segments allows to overcome the problem of weak error correlations.
There are several directions to follow. One is to analyze the continuum limit of the large $N$ designs that result in smooth designs. A second one is to validate our error model against experimental data. A third direction to follow is to study the robust designs for QIP systems that include several or many quantum gates.
Acknowledgements We would like to thank J. Drori and Y. Piasetzky for their assistance in extracting parameters using Lumerical simulations. This work is supported in part by the Israeli Science Foundation Excellence Center, the US-Israel Binational Science Foundation, and the Israel Ministry of Science.
Appendix A Single Qubit Gates
The time evolution of a two-level systems (qubit) is governed by the Schrödinger equation:
$i\partial_{t}\left(\begin{array}[]{l}c_{1}(t)\\ c_{2}(t)\end{array}\right)=\left(\begin{array}[]{cc}-\Delta(t)&\kappa^{*}(t)\\ \kappa(t)&\Delta(t)\end{array}\right)\left(\begin{array}[]{l}c_{1}(t)\\ c_{2}(t)\end{array}\right)\ ,$ (28)
where $c_{1}(t)$ and $c_{2}(t)$ are the probability amplitudes at time $t$ of being in the states $|0\rangle$ and $|1\rangle$, respectively, $\kappa(t)$ is the (complex) Rabi frequency, $\Delta(t)$ is the (real) mode-mismatch, and we set $\hbar=1$. The unitary propagator of such a system is:
$U(t,0)=\mathcal{T}\left[\exp\left[-i\int_{0}^{t}\left(\begin{array}[]{cc}-\Delta\left(t^{\prime}\right)&\kappa^{*}\left(t^{\prime}\right)\\ \kappa\left(t^{\prime}\right)&\Delta\left(t^{\prime}\right)\end{array}\right)\right]dt^{\prime}\right]\ .$ (31)
When $\kappa$ and $\Delta$ are time independent, the propagator simplifies to
$U(t,0)=\exp\left[-it\left(\begin{array}[]{cc}-\Delta&\kappa^{*}\\ \kappa&\Delta\end{array}\right)\right]\ .$ (34)
Physical systems that exhibit this $SU(2)$ dynamics implement one-qubit gates as well as probabilistic two-qubit gates. In silicon photonics, the analog of the time is the propagation axis, and the coupling coefficient $\kappa$ is real. Therefore, for a single segment of length $z$:
$U(\kappa,\Delta,z)=\exp\left[-iz\left(\kappa X-\Delta Z\right)\right]\ ,$ (35)
where $X$ and $Z$ are the Pauli matrices. We can rewrite the $SU(2)$ quantum gate as:
$U(\kappa,\Delta,z)=\exp\left[-iz\Omega_{g}\left(\frac{\kappa}{\Omega_{g}},0,\frac{\Delta}{\Omega_{g}}\right)\cdot\vec{\sigma}\right]$ (36)
$=\cos\left(z\Omega_{g}\right)I-\frac{\kappa}{\Omega_{g}}\sin\left(z\Omega_{g}\right)X-\frac{\Delta}{\Omega_{g}}\sin\left(z\Omega_{g}\right)Z\ ,$ (37)
where $\Omega_{g}\equiv\sqrt{\kappa^{2}+\Delta^{2}}$.
Appendix B Power Loss Simulation Comparison
In Fig. 10 we show a comparison between the power loss simulator (left) to the Fresnel coefficient (right). We consider the FDTD the most accurate simulation method and this is the reason we present this method along with the Fresnel coefficients. The Fresnel coefficient approximation of the power loss compared to the simulation, are related by a global factor of about 5 (see 6 in the main text).
Figure 10: Our power loss simulation using EME (top-left), Fresnel coefficients (top-right), varFDTD (bottom-left) and FDTD (bottom-right). The simulation evaluated the power loss due to a single abrupt transition from width $W_{1}$ (x axis in nm) to width $W_{2}$ (y axis in nm), for Silicon and $H=220nm$.
Appendix C Table of designs
In table 1 we present a list of designs of the Hadamard gate in Lithium Niobate and their fidelity and power loss.
Table 1: Table of Hadamard gate designs in Lithium Niobate. Mat stands for material and PL for Power loss.
[…]

### v2

@@
-Large Segmentation Design of Robust Photonic Quantum Gates
+Robust photonic quantum gates with large number of waveguide segments
 Abstract
-We analyze the fidelity and power loss of photonic quantum gates, implemented by a segmentation of coupled optical waveguides, as functions of the number of segments.
-We show that designs with a large number of segments mitigate effectively the systematic errors for a wide range of error correlation strength and variance.
-We quantify the design roughness effect on the power loss and analyze the performance of various methods for smoothing the configuration of the waveguides.
-We observe an asymptotic scaling behaviour of the gate fidelity and power loss with the number of segments.
+Realizing quantum information processors is challenged by errors and noise across all platforms.
+While composite segmentation schemes have been proposed in many systems, their application to photonic quantum gates in dual-rail configurations has only recently been demonstrated.
+However, prior research has been limited to a small number of segments, full noise correlation, and has overlooked the inherent power loss in such designs.
+Here, we study the fidelity and power loss of composite designs for photonic quantum gates with a high number of segments of varying geometrical widths.
+Using numerical simulations, we analyze the relationship between gate performance and the number of waveguide segments, accounting for statistical error correlations and variances.
+Beyond effectively reducing the errors, an asymptotic scaling pattern of quantum gate fidelity and power loss is observed as the number of segments increases.
+This analysis is examined in Silicon and Lithium Niobate platforms, addressing practical implementation challenges.
+Our findings demonstrate that optimized multi-segment waveguide geometrical designs significantly enhance the robustness and efficiency of photonic quantum gates, paving the way for more reliable quantum information processors.
 I Introduction
-The quest for harnessing the potential of quantum phenomena has been a driving force behind intensive fundamental research and technology efforts, including Quantum Information Processing (QIP) bennett2000quantum .
-Quantum computers hold the promise for an exponential speedup in solving certain complex computational problems and a capability to decrypt classical encryption protocols in real-time, thereby opening new frontiers in computing and cryptography.
-Despite substantial theoretical advancements in quantum algorithms QuantumAlgorithms1 ; QuantumAlgorithmSurvey , the practical realization of a quantum information processor remains a formidable challenge.
-One of the paramount obstacles in achieving practical quantum computation lies in the noise and errors that affect the performance of the fundamental building blocks of quantum algorithms: the qubit’s state preparation, the quantum gates, and the measurement process Nielsen2000 .
-These errors and noise have various sources, such as fabrication imperfections Kyoseva2019 , inaccurate coupling strengths and the interaction with the classical environment noiseoneoverf .
-In response to these challenges, two primary routes have been taken: error-correcting codes QEC1 and error mitigation schemes NoiseMitigation2 ; QNoiseMitigation1 .
-Error-correcting codes construct logical qubits using additional entangled physical qubits and can significantly outperform their physical counterparts.
-Notable among these codes are stabilizer codes, exemplified by the surface code Fowler2012 , which exhibits a relative tolerance to local errors.
-Fault-tolerant quantum computation using error correcting codes requires the fidelity of the physical quantum gates to be above a certain threshold, making the reduction of the physical errors to be of utmost importance NoiseThreholdQC ; QuantumErrorCorrection .
-Error mitigation schemes seek to reduce errors using sophisticated implementation and techniques, for a review see QNoiseMitigation1 ; NoiseMitigation2 .
-One such framework is the segmented composite pulse design, which was originally used in atomic physics and nuclear magnetic resonance, employing a combination of constant pulses to minimize errors during the system evolution CPNMR ; CPNMR2 ; CPAtomicPhysics ; CPAtomicPhysics2 .
-This error mitigation technique holds a promise for improving the performance of complex quantum algorithms, and has a broad applicability in diverse quantum computing realizations of quantum information processing devices.
-In SmoothCP smooth designs were proposed in a general composite pulses scheme, and HUSAIN2013145 suggested symmetric and antisymmetric composite pulses with a large number of pulses in NMR.
-It has been recently shown in Kaplan2023 that this method can be implemented in Silicon photonics KLM for designing robust high fidelity quantum gates by varying the waveguide geometry, and specifically its cross-section.
-The focus in Kaplan2023 has been on having a small number of segments, under the assumption of fully correlated noise of the different waveguides.
-It was proven in YaronIdoNoiseThreshold that there is a correlation strength threshold for the segmented design to be an effective error mitigating scheme.
-The aim of the current work is to analyze the composite pulse scheme for optimizing robust designs with a large number of segments.
-We consider both the fidelity of the quantum gates as well as the power loss of the designs for varying systematic noise correlations.
-Applied to the photonic dual-rail realization, our methods yield robust high-fidelity unitary designs for single-qubit and hence also for the probablistic double-qubit gates such as CZ and CNOT Knill2001 .
-We consider different manufacturing configurations in Silicon and Lithium Niobate, as Silicon is widely used in this field, while the Lithium Niobate is considered to be a promising material for designing active QIPs.
-Our main contributions are :
-  1. 1.
-    Enhanced Quantum Gate Robustness: We demonstrate numerically that increasing the number of composite segments $N$ results in a substantial improvement in the quantum gate mean fidelity $\mathbf{E}\left[F\right]$, and a reduction in its standard deviation $\sigma_{F}\left[F\right]$ with a certain scaling with $N$.
-We further show that $N\sim O(10)$ is effective for most practical cases.
-  2. 2.
-    Power Loss Analysis: We quantify the impact of increasing $N$ on the power loss in the quantum gate due to the back-scattering effects in the segments.
-We argue analytically and show numerically that the power loss grows linearly with the number of segments, which makes it necessary to apply smoothing methods to the large $N$ designs.
-  3. 3.
-    Correlation Strength: We study the impact of error correlation strength between the different segments on the gate fidelity.
-We show that enhancing the correlation strength among the errors in different segments leads to a significant increase in the expected fidelity of the quantum gate, hence providing valuable insights to strategies for choosing the optimal number of segments given a correlation strength.
-For a given required gate robustness, larger correlation strength implies that less segments are needed.
-  4. 4.
-    Adiabatic and Fresnel Designs: We introduce smooth segmented gates designs based on the principles of the Adiabatic and Fresnel scattering, that exhibit a reduced power loss while keeping the same fidelity performance.
-This offers a complementary tool to the traditional segmentation method.
-The paper is organized as follows: In Section II, we briefly review the composite pulse design of quantum gates and the error model in the optical waveguides framework.
-In Section III we present the ingredients of the cost function optimization framework including the optimization range, the quantum gate fidelity and the smooth design conditions for minimizing the power loss.
-In Section IV we show the optimization results for the gates fidelity and power loss for a varying number of segments and for different smooth designs.
-Section V is devoted to a discussion and conclusions.
-Additional details are provided in four appendices.
-II Segmented Composite Pulses
-Figure 1: We divide each of the two waveguides to $N$ segments.
+The realization of a practical quantum information processor remains a great challenge, as errors and noise inevitably arise during the implementation of its core elements.
+These challenges are universal across various physical platforms for quantum information processing bennett2000quantum ; Nielsen2000 , including superconducting circuits Kjaergaard2020 , trapped ions Bruzewicz2019 , and photonic systems Flamini2018 ; giordani2023integrated .
+Among these, integrated photonics, particularly dual-rail waveguide platforms, has emerged as a promising approach due to its scalability and compatibility with existing fabrication technologies.
+In dual-rail photonic systems, errors primarily stem from fabrication imperfections Kyoseva2019 , inaccurate coupling strengths, and environmental interactions noiseoneoverf , which compromise the fidelity of the quantum state preparation and quantum gates, the fundamental building blocks of any quantum algorithm Nielsen2000 .
+For example, deviations in waveguide geometries from the intended design, such as variations in waveguide width, can significantly impact the performance of dual-rail single-photon and two-photon gates.
+These deviations induce Pauli errors, which in turn lead to a reduction in gate fidelity melati2014real .
+Various strategies have been developed to address these challenges QNoiseMitigation1 ; NoiseMitigation2 .
+Error correction codes, for example, construct logical qubits from entangled physical qubits to reduce errors QEC1 .
+However, these methods impose stringent fidelity thresholds on the underlying physical gates, requiring a fidelity error of less than $1\%$ in each of the physical photonic gates NoiseThreholdQC ; QuantumErrorCorrection ; shi2022high ; noiri2022fast .
+One promising error mitigation approach, which reduces physical errors through advanced design and implementation techniques, is the composite design method, originally developed in atomic physics and nuclear magnetic resonance (NMR) CPNMR ; CPNMR2 ; CPAtomicPhysics ; CPAtomicPhysics2 .
+This method employs systematic parameter variations to counteract errors during system evolution.
+In NMR, large number of pulses were suggested, using symmetric and antisymmetric composite pulses designs HUSAIN2013145 , and a smooth composite pulses scheme have been proposed in Ref.
+SmoothCP .
+Recent work has adapted this approach to integrated photonics, demonstrating the potential for high-fidelity dual-rail quantum gates by segmenting waveguides and varying their geometries Kyoseva2019 ; Kaplan2023 .
+However, prior studies primarily considered designs with a small number of segments and assumed fully correlated fabrication errors Kaplan2023 .
+Subsequent findings indicate that effective error mitigation requires the correlation strength to remain below a specific threshold YaronIdoNoiseThreshold .
+Moreover, earlier analyses did not evaluate the power loss associated with segmentation, an essential factor for practical implementation.
+Figure 1: Comparison of composite segment design with variable number of segments.
+(a) Two coupled waveguides, with widths denoted as $W_{0}$ and $W_{1}$, a gap between them and a defined length.
+(b) Optical waveguide gates in dual-rail encoding for $X$ (Pauli), gate, along with its graphical representation on the Bloch sphere.
+(c) Fabrication noise model with spatial correlation between segments, for a small and large number of segments.
+(d) Comparison of Fidelity and Power Loss between a small and large number of segments.
+Increasing either the number of segments or the noise spatial correlation improves Fidelity.
+Our proposed smooth design significantly reduces power loss, in addition to increasing Fidelity.
+Here, we study the fidelity and power loss of photonic quantum gates composed of increased number of waveguides’ segmentations.
+We find that as the number of segments increases, the errors associated with variations in waveguide geometrical widths are significantly mitigated, leading to substantially higher gate fidelity.
+Using numerical simulations on Silicon-on-Insulator (SOI) and Thin-Film Lithium Niobate (TFLN) platforms, we demonstrate that increasing the number of waveguide segments $N$, not only improves the average quantum gate fidelity $\mathbf{E}\left[F\right]$ but also reduces its standard deviation $\sigma_{F}\left[F\right]$, with a specific scaling with $N$.
+Furthermore, we analyze the impact of segmentation on power loss due to back-scattering effects.
+Our analytical and numerical results show that the power loss increases linearly with the number of segments, underscoring the necessity of employing smoothing techniques for designs with large $N$.
+To address this, we propose smooth segmented photonic quantum gate designs based on adiabatic and Fresnel reflection principles.
+These designs effectively reduce power loss while maintaining high fidelity performance.
+Finally, we conclude that a segmentation number $N\sim O(10)$ is sufficient for most practical implementations, providing a balance between fidelity enhancement and manageable power loss.
+This work advances the development of practical quantum dual-rail gates for photonic integrated platforms, and quantum information processors in integrated photonics.
+The paper is organized as follows: Section II introduces the theoretical background, including dual-rail encoding within the framework of integrated optics, as well as the necessary background on composite schemes.
+In Section III, we formulate the optimization problem and define the cost function, incorporating the optimization range, quantum gate fidelity, and smooth design conditions aimed at minimizing power loss.
+Section IV presents the optimization results for gate fidelity and power loss across varying numbers of segments and different smooth design configurations.
+Finally, Section V provides a summary and outlook.
+Additional technical details are included in four appendices.
+II Theoretical background
+II.1 Dual-rail realization for of two-mode dynamics
+In the integrated dual-rail realization, two closely spaced optical waveguides form a coupled system that emulates the dynamics of a quantum two-level system AmnonYariv .
+This system, governed by coupled-mode theory (CMT), is described by a unitary propagator that captures its evolution over the propagation length:
+$i\partial_{z}\left(\begin{array}[]{l}E_{0}(z)\\ E_{1}(z)\end{array}\right)=\left(\begin{array}[]{cc}-\Delta(z)&\kappa^{*}(z)\\ \kappa(z)&\Delta(z)\end{array}\right)\left(\begin{array}[]{l}E_{0}(z)\\ E_{1}(z)\end{array}\right)\ ,$ (7)
+where $E_{0}(z)$ and $E_{1}(z)$ are the electrical fields at the of being in states $|0\rangle$ and $|1\rangle$, the mode mismatch $\Delta(z)$ arises from differences in the effective refractive indices, $n_{\text{eff},1}$ and $n_{\text{eff},2}$ of the waveguides:
+$\Delta(z)=\frac{\beta_{1}(z)-\beta_{2}(z)}{2}\equiv\frac{2\pi}{\lambda}\left(n_{{eff}_{1}}(z)-n_{{eff}_{2}}(z)\right)\ ,$ (8)
+where $\beta_{1}$ and $\beta_{2}$ are the propagation constants of the two waveguides.
+The coupling coefficient $\kappa$, resulting from the evanescent field overlap between the waveguide modes, is defined as:
+$\kappa(z)=\frac{c}{4\pi\lambda}\iint$ $\left[\epsilon(x,y,z)-\epsilon^{(1)}(x,y,z)\right]\cdot$
+[…]
+where $(0)$ and $(1)$ refer to the two waveguides, $(x,y)$ are the transverse directions, $z$ denotes the propagation direction, $\vec{E}_{\perp}^{(0)}(x,y,z)$ and $\vec{E}_{\perp}^{(1)}(x,y,z)$ are the transverse electric field modes for each of the waveguides, and $\epsilon(x,y,z)$ and $\epsilon^{(1)}(x,y,z)$ are the permittivity distributions of the coupled waveguides system (namely, the profile of the two waveguides together) and of waveguide $(1)$ alone, respectively AmnonYariv .
+$c$ is the speed of light and $\lambda$ is the wavelength.
+A diagram of two coupled waveguides is shown in Fig.
+1 a.
+The unitary propagator of such a system is:
+$U(z,0)=\mathcal{Z}\left[\exp\left[-i\int_{0}^{z}\left(\begin{array}[]{cc}-\Delta\left(z^{\prime}\right)&\kappa^{*}\left(z^{\prime}\right)\\ \kappa\left(z^{\prime}\right)&\Delta\left(z^{\prime}\right)\end{array}\right)\right]dz^{\prime}\right]\ ,$ (12)
+where $\mathcal{Z}$ denotes z-ordering. In cases where $\kappa$ and $\Delta$ are independent of z, the propagator simplifies to:
+$U(z,0)=\exp\left[-iz\left(\begin{array}[]{cc}-\Delta&\kappa^{*}\\ \kappa&\Delta\end{array}\right)\right]$ (15)
+Alternatively, using the well-known relation $e^{-i\alpha\vec{\sigma}\cdot\hat{n}}=cos(\alpha)\mathbf{I}+\vec{\sigma}\cdot\hat{n}sin(\alpha)$, for a unit vector $\hat{n}$, and $\vec{\sigma}=(\mathbf{X},\mathbf{Y},\mathbf{Z})$ a vector of Pauli matrices, the unitary operator can be written as:
+$U\left(\kappa,\Delta,z\right)=\cos\left(\Omega_{g}z\right)\mathbf{I}-i\frac{\kappa}{\Omega_{g}}\sin\left(\Omega_{g}z\right)\mathbf{X}$
+$+i\frac{\Delta}{\Omega_{g}}\sin\left(\Omega_{g}z\right)\mathbf{Z}\ ,$ (16)
+while we use use the transfer frequency determined by the effective coupling parameter $\Omega_{g}=\sqrt{\kappa^{2}+\Delta^{2}}$.
+The dual-rail encoding in two coupled optical waveguides forms a quantum two-level system, where light propagation emulates $SU(2)$ dynamics, and is elegantly represented on the Bloch sphere (Fig.
+1 (b)), where the evolution of light in the coupled waveguides mirrors Rabi oscillations in atomic physics.
+For symmetric configurations ($\Delta=0$), the system achieves Complete Power Transfer (CPT), between the two levels.
+This $SU(2)$ framework underpins the implementation of integrated photonic quantum gates and sets the stage for advanced multi-qubit operations via concatenated or segmented couplers.
+Realizing two-qubit gates, such as the Controlled-NOT (CNOT) or Controlled-Z (CZ) gates, typically requires additional ancilla modes, measurements, and linear optical elements, and these gates often exhibit a non-deterministic nature KLM .
+A well-known example is the CNOT gate, which can be implemented using beam splitters, phase shifters, and single-photon detectors Ralph2002 .
+By carefully arranging these components, conditional operations are performed, where the state of one qubit (the control qubit) governs the state of another (the target qubit).
+The universality of quantum computing in the dual-rail encoding scheme is achieved through the combination of single-qubit and two-qubit gates.
+This universal gate set enables the construction of arbitrary quantum circuits, paving the way for the implementation of complex quantum algorithms.
+II.2 Coherent Noise Analysis
+Figure 2: $N$ segments composite configuration.
+We divide each of the two waveguides to $N$ segments.
 Each segment $i$ is characterized by its coupling $\kappa_{i}$, mode-mismatch coefficient $\Delta_{i}$ and length $z_{i}$.
-The interpolation map for $\kappa$ and $\Delta$ as a function of the waveguides widths $W_{1},W_{2}$, in $nm$ units.
-These results are extracted from a Lumerical FDTD Eigensolver simulation, for Silicon on insulator waveguides platform with a height of $220nm$ and a gap of $1\mu m$.
-One of the waveguides was illuminated with $1.55\mu m$ light source.
-II.1 Robust Composite Quantum Gates
-The presence of noise within the physical two-level system leads to deviations from the desired quantum gate operation.
+The interpolation map for $\kappa$ and $\Delta$ as a function of the waveguides widths $W_{0},W_{1}$, in $nm$ units.
+These results are extracted from a Lumerical FDTD Eigensolver simulation, with geometry resolution of $5nm$, and a bilinear interpolation to evaluate the values in-between.
+Coherent noise in photonic quantum gates, caused by fabrication imperfections, introduces systematic errors in the form of random unitary rotations.
+In the photonic regime, these errors are primarily attributed to variations in the waveguide widths during the fabrication process, which significantly impact the coupling coefficient $\kappa$ and mode mismatch $\Delta$.
+Additionally, due to the nature of errors arising during the fabrication of coupled waveguides, these errors are fully correlated between the two coupled waveguides, further compounding their effect on gate fidelity ExperimentCP .
+Such systematic deviations degrade gate performance and accumulate in multi-qubit operations, necessitating mitigation to meet the stringent thresholds required for quantum error correction and measurement-based quantum computation Nielsen2000 .
 To quantify the extent to which the noisy gate deviates from the ideal gate, we define the gate fidelity $F$ by the trace norm:
 $F\left(U_{\text{ideal}},U(\boldsymbol{\epsilon})\right)=\frac{1}{2}\left|\text{Tr}\left(U_{\text{ideal}}^{\dagger}U(\boldsymbol{\epsilon})\right)\right|\ ,$ (17)
-where $U_{\text{ideal}}$ represents the desired ideal unitary gate, and $U(\boldsymbol{\epsilon})$ represents its physical realization, which depends on a set of $m$ jointly distributed random errors denoted as $\boldsymbol{\epsilon}=\{\epsilon^{a}\}_{a=1}^{m}$.
+where $U_{\text{ideal}}$ represents the desired ideal unitary gate, and $U(\boldsymbol{\epsilon})$ represents its physical realization, which depends on a set of $m$ jointly distributed random errors denoted as $\boldsymbol{\epsilon}=\{\epsilon_{a}\}_{a=1}^{m}$.
 The fidelity is a random function that takes values in the interval range $[0,1]$, where zero and one correspond to the absence of errors and to the maximal deviation from the desired quantum gate operation, respectively.
-The primary objective of our analysis is to maximize the mean fidelity, denoted as $\mathbf{E}\left[F\right]$, and to minimize its standard deviation, denoted as $\sigma[F]$:
+The primary objective of our analysis in the context of error mitigation, is to maximize the mean fidelity, denoted as $\mathbf{E}\left[F\right]$, and to minimize its standard deviation, denoted as $\sigma[F]$:
 $\mathbf{E}\left[F\right]$ $=$ $\mathbb{E}_{\boldsymbol{\epsilon}}\left[F\left(U_{\text{ideal}},U(\boldsymbol{\epsilon})\right)\right]\ ,$
 $\sigma[F]$ $=$ $\sqrt{\mathbb{E}_{\boldsymbol{\epsilon}}\left[F\left(U_{\text{ideal}},U(\boldsymbol{\epsilon})\right)^{2}\right]-\mathbf{E}\left[F\right]^{2}}\ .$ (18)
-The composite pulses is a framework for error mitigation where $U(\boldsymbol{\epsilon})$ is realized as a product of $N$ physical unitary gates:
-$U_{cp}(\vec{\boldsymbol{\epsilon}})=\prod_{i=1}^{N}U_{i}(\boldsymbol{\epsilon}_{i})\ ,$ (3)
-where $\vec{\boldsymbol{\epsilon}}=(\boldsymbol{\epsilon}_{1},...,\boldsymbol{\epsilon}_{N})$, $\boldsymbol{\epsilon}_{i}=\{\epsilon^{a}_{i}\}_{a=1}^{m}$, and $U_{i}$ are constructed to achieve higher fidelity.
-We will consider this error mitigation scheme in the framework of dual-rail waveguides in integrated photonics.
-II.2 Coupled Optical Waveguides
-Consider the following physical photonic realization of single-qubit quantum gates, which is also the basis for constructing probabilistic two-qubit gates AmnonYariv .
-A coupled system of two waveguides can be described by two parameters: the coupling coefficient $\kappa$ and the mode-mismatch coefficient $\Delta$ (known also as the detuning coefficient).
-The coupling coefficient is given by the integral over the spatial modes-overlap:
-[…]
-where $(1)$ and $(2)$ refer to the two waveguides in the dual-rail representation, $(x,y)$ are the transverse directions, $z$ denotes the propagation direction and $\epsilon(x,y)$ and $\epsilon^{(2)}(x,y)$ are the permittivity distributions of the coupled waveguides system and of waveguide $(2)$ alone, respectively.
-The mode-mismatch coefficient is determined by the difference between the propagation constants of the two waveguides:
-$\Delta=\frac{\beta_{1}-\beta_{2}}{2}\equiv\frac{2\pi}{\lambda}\left(n_{{eff}_{1}}-n_{{eff}_{2}}\right)\ ,$ (5)
-where $n_{{eff}_{i}}$ are the effective refraction indices in the first and second waveguides, respectively, and depend on the geometry of the waveguides.
-A simple solution for the propagation differential equation is constant coupling and mode-mismatch, which we refer to as one segment.
-It is given by (see appendix A):
-$U\left(\kappa,\Delta,z\right)=\cos\left(\Omega_{g}z\right)\mathbf{I}-i\frac{\kappa}{\Omega_{g}}\sin\left(\Omega_{g}z\right)\mathbf{X}$
-$+i\frac{\Delta}{\Omega_{g}}\sin\left(\Omega_{g}z\right)\mathbf{Z}\ ,$ (6)
-where $\Omega_{g}\equiv\sqrt{\Delta^{2}+\kappa^{2}}$, and $\mathbf{I},\mathbf{X},\mathbf{Z}$ are the identity matrix and Pauli matrices, respectively.
-II.3 Segmentation
-We will analyze quantum gates designs that consist of $N$ segments (see Fig.
-1).
-Supported by experimental evidence ExperimentCP , we consider the case where in each segment, the noise affects both waveguides identically (full correlation), i.e.
-$\epsilon^{(1)}_{k}=\epsilon^{(2)}_{k},~{}k=1,...,N$.
-This implies a possible significant error in the coupling coefficients $\kappa_{k}$, but affects less the mode-mismatch coefficients $\Delta_{k}$.
-As to the correlation between the errors of the different segments, we will initially assume a full correlation between them, and will later relax this condition and analyze the case of a general correlation matrix.
-The $N$-segmented gate reads:
-$U^{(N)}=\prod_{k=1}^{N}U(\kappa_{k},\Delta_{k},z_{k},\epsilon_{k})\ ,$ (7)
-where $z_{k},\kappa_{k},\Delta_{k},\epsilon_{k}$ are the length of the $k^{th}$-segment, its coupling, its mode-mismatch coefficient and its noise, respectively.
-II.4 Power Loss
-Implementation of the composite pulse error mitigation scheme in Silicon photonics or in any other physical platform is accompanied by power loss due to back-scattering.
-This back-scattering is a direct consequence of the non-smooth profile of the segmented waveguide’s width, and captures the total power loss (see Appendix B).The gate reflection power loss is total power transmission coefficient $P_{t}$:
+Implementation of the composite pulse error mitigation scheme in integrated photonics, or in any other physical platform, is accompanied by power loss due to several reasons such as absorption or scattering.
+The back-scattering is a direct consequence of the non-smooth profile of the segmented waveguides width profiles, and captures the total power loss (see Appendix B).
+The gate reflection power loss is total power transmission coefficient $P_{t}$:
 $P_{t}\equiv T=\prod_{k=1}^{N-1}T_{k}(\kappa_{k},\Delta_{k},z_{k},\epsilon_{k})\ ,$ (19)
 where $T_{k}$ is the power transmission coefficient from the segment $k$ to the segment $k+1$.
-Specifically, in photonics, we can estimate the back-scattering by the Fresnel equation that governs the reflection and transmission coefficients of a light transmitted between two different materials with refractive indices $n_{1}$ and $n_{2}$ (we consider the intensity ):
-$R=\left(\frac{n_{1}cos\theta_{i}-n_{2}cos\theta_{t}}{n_{1}cos\theta_{i}+n_{2}cos\theta_{t}}\right)^{2}\ ,$ (9)
-where we chose TE polarization and $\theta_{i}$ and $\theta_{t}$ are the incident and transmitted light angle, respectively.
-In guided propagation, the refractive index is replaced by $n_{eff}$, which is calculated from the overlap integral of the spatial cross-section field profile with the cross-section refractive index.
-For our analysis, focusing on the TE polarization (the basic mode) and assuming small angles related to the $z$ axis, we evaluate the total total reflection coefficient as:
-$R=1-T=1-\prod_{i=1}^{N-1}\left(1-\left(\frac{n_{eff_{i+1}}-n_{eff_{i}}}{n_{eff_{i+1}}+n_{eff_{i}}}\right)^{2}\right)\ .$ (10)
-III Optimization Framework
-Our optimization search method is subject to several constraints.
-While our goal is to obtain high fidelity quantum gates, we aim to achieve it with as smooth as possible design in order to minimize the power loss.
-As discussed, abrupt changes in the waveguide’s width increase the power loss due to back scattering, and also give rise to higher spatial modes in the waveguide that distort the signal, hence changing the qubit state.
+Due to the changes in the effective dispersion coefficients between the segments of the waveguides, we estimate the back-scattering by the Fresnel equation, with effective refractive indices of the two consecutive segments:
+$T_{k}=1-\left(\frac{n_{\text{eff},k}-n_{\text{eff},k+1}}{n_{\text{eff},k}+n_{\text{eff},k+1}}\right)^{2}\ ,$ (20)
+The total reflection coefficient is given by:
+$R=1-P_{t}\ .$ (21)
+II.3 The Composite Segmentation Method
+The composite pulses method is well-known for its ability to mitigate errors by decomposing a single gate into a sequence of smaller unitary operations.
+By strategically leveraging the correlation between error factors, this sequence can be meticulously designed to cancel out errors effectively.
+In the dual-rail implementation, as demonstrated in Ref.
+Kaplan2023 , a two-waveguide coupler is divided into multiple sections with varying widths and shorter propagation lengths.
+We define the gate $U(\boldsymbol{\epsilon})$ as a sequence of $N$ segments:
+$U_{cp}(\vec{\boldsymbol{\epsilon}})=\prod_{i=1}^{N}U_{i}(\boldsymbol{\epsilon}_{i})\ ,$ (22)
+Where $\boldsymbol{\epsilon}_{i}$ are the error parameters (in our context, in the width of the waveguides), a single one for each gate.
+The $U_{i}$ are characterized by a its own $\kappa_{i},\Delta_{i},z_{i}$ parameters (coupling coefficient, mode mismatch, and length, respectively).
+Therefore, designing a robust unitary gate $U_{cp}$ is actually finding the set of geometrical parameters (width, length, refractive index etc.), that determines $\{\kappa_{i},\Delta_{i},z_{i}\}_{i=1}^{3}$ (total of $3N$ degrees of freedom) such that the gate will be more robust to the presence of noise, namely, the fidelity between $U_{cp}$ and $U$ will be as close as possible to 1.
+III High fidelity quantum gates using Large N segmentation
+In this study, we focus on dual-rail waveguide platforms, presenting a detailed study of segmented and smooth waveguide designs of robust, high-performance quantum gates designs.
+Fig.
+1 provides an overview of the work, including the key results from this paper: showcasing the relationship between fidelity and power loss against the number of segments (see Fig.
+1c), as well as the impact of smoothing and noise correlation (see Fig.
+1d).
+Our aim is to obtain high-fidelity quantum gates with as smooth as possible design that will minimize the power loss, for any arbitrary number of segments $N$.
+As discussed, abrupt changes in the waveguides widths increase the power loss due to back scattering, and also give rise to higher spatial modes in the waveguide that distort the signal, hence changing the qubit state.
 Moreover, waveguide designs must be physically implementable subject to the manufacturing resolution.
 We divide each waveguide into a sequence of $N$ small segments, each of length $z_{i}$ as shown in Fig.
-1.
-The two waveguides are separated by a gap (defined as the distance between the centers of the waveguides) and have a height $h$.
-Our optimization uses the Lumerical FDTD results as a look-up table to inverse-map the geometry to the coupling coefficient and the mode-mismatch parameters.
-III.1 Fidelity
-We will assume that the errors in the two waveguides widths $W_{1}$ and $W_{2}$ are fully correlated and can be sampled from a Normal distribution with zero mean and a standard deviation $\sigma$, $D\sim\mathcal{N}(0,\sigma)$.
+2.
+The two waveguides are separated by a gap, defined as the distance between the centers of the waveguides, and have a height $h$.
+The optimization process involves three stages.
+Firstly, we construct a look-up table with the Lumerical FDTD simulator to map geometries to the coupling coefficient and mode-mismatch parameters inversely.
+Next, we establish the optimization cost function, including regularizes.
+Finally, we filter-out devices that are not physically realizable.
+The subsequent subsections will cover these stages in depth.
+III.1 The Fidelity Term
+As discussed, in our analysis we assume that the errors in the two waveguides widths $W_{0}$ and $W_{1}$ are fully correlated, and can be sampled from a Normal distribution with zero mean and a standard deviation $\sigma$, $D\sim\mathcal{N}(0,\sigma)$.
 For our analysis, we consider the dimensionless parameter $\sigma_{\%}\equiv 100\frac{\sigma}{W_{mean}}$, where
 $W_{mean}\equiv\frac{1}{2N}\sum_{i}\left(W_{i}^{(0)}+W_{i}^{(1)}\right)\ ,$ (23)
-and we focus on the range $1\%-10\%$, which is approximately around $10_{nm}-50_{nm}$. The fidelity loss term reads:
-$\mathcal{L}_{F}=1-E_{\delta W\sim D}\left[\frac{1}{2}Tr\left(U_{ideal}^{\dagger}U\right)\right]\ ,$ (12)
+and we focus on the range $1\%-10\%$, which is approximately $10_{nm}-50_{nm}$. The fidelity loss term reads:
+$\mathcal{L}_{F}=1-E_{\delta W\sim D}\Big{[}F\left(U_{\text{ideal}},U\right)\Big{]}\ ,$ (24)
 where $U_{ideal}$ is the ideal quantum gate and $U$ is its physical realization.
-III.2 The Range of Width
+III.2 The Range of Width and Length
 We add a regularization in order to restrict the width of the waveguides to be in a range between the minimum and the maximum of an allowed region:
@@
 where $\lambda_{W}$ is a hyperparameter for the minimal and maximal widths.
-We define the optimization domain $\Omega=(\Omega_{dz},\Omega_{w})$ as the set of all experimentally realizable waveguide widths.
+We define the optimization domain $\Omega=(\Omega_{dz},\Omega_{w})$, as the set of all experimentally realizable waveguide widths.
 The number $N$ of segments in the design varies between 1 and 200, and we denote by $L$ the total length of the design.
@@
 The waveguide width initialization is sampled uniformly in the range between the maximum and minimum of the feasible domain.
-Additionally, strong regularization is activated once the design is out of the optimization domain $\Omega$.
-It is also possible to add another regularization for the total design length $L$, but in our analysis we do not impose this restriction.
-Almost all the design lengths are between $10\mu m$ and $100\mu m$.
+Furthermore, strong regularization is activated once the design is out of the optimization domain $\Omega$.
+To avoid overly short segments that could violate our assumptions and necessitate high fabrication resolution, we restricted our focus to devices with sufficiently large segment lengths.
+Specifically, we ensured that the dimensionless parameter characterizing propagation, defined as the ratio of each segment’s length to the wavelength in the material, remains sufficiently large.
+While it is also possible to impose an additional regularization on the total design length $L$, this restriction is not applied in our analysis.
+Instead, filtered-out devices with a minimum segment length less than $1\mu m$.
+Our results show that nearly all designs naturally exhibit total lengths in the range of $10\mu m$ to $100\mu m$ (see the table of our designs 1 in the Appendix).
 III.3 Smooth Design Conditions
-III.3.1 Baseline optimization
-We define the baseline optimization as a search for the highest fidelity designs even with having abrupt changes between segments cross-section widths.
-This exhibits scattering and reflections, higher spatial modes generations and power loss, and requires a very high precision for manufacturing.
-Designs will be compared to the baseline ones.
-III.3.2 Minimizing Fresnel Reflection
-In order to reduce the power loss, we add a regularization term which reduces the back reflection coefficient.
+We define the baseline optimization, as a search for the highest fidelity designs, without taking into account any smoothness requirement.
+This lead to designs with high back-scattering, and power loss.
+The alternative optimized designs will be evaluated in comparison to these baseline configurations.
+III.3.1 Minimizing Fresnel Reflection
+In order to reduce the power loss, we add a regularization term, which reduces the back reflection coefficient.
 Thus, among the robust high-fidelity design parameter space, we search for the designs with the minimal power loss.
@@
 $\mathcal{L}=\mathcal{L}_{F}+L_{\Omega}+\lambda_{R}R\ ,$ (26)
-where $\mathcal{L}_{F},L_{\Omega},R$ are given by (12), (13) and (10), respectively.
-III.3.3 Imposing An Adiabaticity Condition
+where $\mathcal{L}_{F},L_{\Omega},R$ are given by Eq. 24, 25 and 21, respectively.
+III.3.2 Imposing An Adiabaticity Condition
 The adiabatic theorem guarantees, under certain conditions, a smooth transition between the ground states of two different Hamiltonians adiabatic .
@@
 $\mathcal{L}=\mathcal{L}_{F}+L_{\Omega}+\lambda_{A}A\ .$ (30)
-Figure 2: Example of two 100-segments designs of $\sqrt{X}$ gate in Silicon, with width noise standard deviation of $\sigma=30$nm ($\sigma_{\%}\approx 5\%$).
-Top: the Fresnel-based design with expectation fidelity of $0.91$ and reflection coefficient of ±$\pm 10^{-3}$.
-Bottom: the baseline with expectation fidelity of $0.9$ and reflection coefficient of ±$\pm 10^{-2}$.
-The roughness standard deviation (width change along the design) is $23$nm and $76$nm, respectively.
-III.4 Error Correlations
-The errors at different segments may be partially correlated.
-In order to quantify the effect on the design performance of different strength of correlations between the errors of different segments we introduce the error correlation matrix between the $k,l\in\{1,...N\}$ segments:
+III.4 Segments Error Correlation Model
+We also account for cases with partial correlation, where errors in different segments may exhibit varying degrees of dependency.
+To assess the impact of correlation strength on design performance, we introduce an error correlation matrix to model the relationships between errors across $k,l\in\{1,...N\}$ segments:
 $\rho_{kl}(\mu)=\begin{cases}1&k=l\\ \frac{1}{\left(\left|k-l\right|+1/2\right)^{2\mu}+1}&k\neq l\ .\end{cases}$ (31)
 The dimensionless parameter $\mu$ controls the strength of the correlation: $\mu=0$ implies strong correlations between segments while $\mu\rightarrow\infty$ implies no correlation.
-The results of our correlations simulation will be discussed in the the simulation section (see IV.3).
-III.5 Waveguide Material
-We optimize the design for two types of materials: Silicon, which is widely used in photonic applications and has a relatively high refractive index (about 3.48), and LiNbO3 which is considered to be a promising candidate for photonic applications due to its outstanding optical properties, including a wide transparency window $(0.35\mu m-5\mu m)$, large optical nonlinearity and electro-optic coefficient, and strong acousto-optic effect.
-The Lithium Niobate has negative uniaxial birefringence, but we assume that only the first $TE_{01}$ mode is occupied.
+The results of our correlations simulation will be discussed in the simulation section (see sub-section IV.3).
+We optimize the design for two materials: Silicon and $\text{LiNbO}_{3}$ (Lithium Niobate).
+Silicon, widely used in photonic applications, is notable for its high refractive index of approximately 3.48.
+$\text{LiNbO}_{3}$, on the other hand, is a promising candidate due to its exceptional optical properties, including a wide transparency window ($0.35\mu m$ to $5\mu m$), large optical nonlinearity, high electro-optic coefficient, and strong acousto-optic effect.
+We assume in our analysis that only the fundamental $TE_{01}$ mode is occupied.
 For both materials we use a gap of $1.0~{}\mu m$ and the width of each waveguide is varied between $500~{}nm$ to $850~{}nm$.
-We further assume manufacture resolution (smallest resolution in the optimization) to be $20~{}nm$, waveguide height of $220~{}nm$, etch of $150~{}nm$ and wavelength of $1.55~{}\mu m$.
+We further assume the manufacture resolution (smallest resolution in the optimization) to be $20~{}nm$, waveguide height of $220~{}nm$, etch of $150~{}nm$ and wavelength of $1.55~{}\mu m$.
 To evaluate the mode-mismatch and coupling coefficients of the design we use Lumerical Eigensolver simulation (as shown in Fig.
-1).
-We simulate different design geometries and build a look-up table to express the mode-mismatch coefficient $\Delta$ and the coupling coefficient $\kappa$ in terms of the geometrical waveguide parameters.
-The resolution of the geometry is $5~{}nm$, and we use a bi-linear interpolation to evaluate the values in-between.
+2).
 IV Optimization Results
@@
 We use the PyTorch Package, with Adam optimizer, CosineAnnealing LR scheduler, with an initial learning rate of 0.1, batch size of 1024 samples, and the number of epochs was set to be around 5000.
-The device length is initialized randomly between $25~{}\mu m$ to $100~{}\mu m$.
+The length of the device is randomly initialized between $25~{}\mu m$ and $100~{}\mu m$.
 In our analysis we focus on the Pauli $X$ (without mode-mismatch) and the Hadamard $H$ (with mode-mismatch) single-qubit quantum gates.
-We found that other single-qubit gates show a similar behaviour.
-In Fig.
-2 we see examples of composite designs (baseline and Fresnel based).
-Diagram of the trajectories on the Bloch sphere of designs for different optimization methods are shown in Fig.
-3.
-Figure 3: Bloch Sphere diagram of two optimization results.
-Left - Hadamard gate with 10 segments, initial state is $|0\rangle$ and final state is |+ket|$\rangle$| + ⟩.
-Right - Pauli X gate with 50 segments, initial state is $|+\rangle_{y}$ and final state is $|-\rangle_{y}$.
+In this study, we observed that the scaling behavior of the two materials is quite similar, as well as the gate types.
+Consequently, we decided to concentrate on presenting our analysis for only one material, and one gate at a time.
+To visually represent the geometric design achieved through the optimization process, refer to Fig. 6. For a depiction of the trajectories on the Bloch sphere, please see Appendix A.
 IV.1 Mean Fidelity and Variance
@@
 3 shows the expectation value and standard deviation of the fidelity for different numbers of segments taken in the range 1 to 200.
-To clean the optimization numerical noise, we avergated over 20 optimizations (for each number of segments).
-As can be clearly seen from the plot, increasing the number of segments leads to an oncrease in the gate fidelity, and the most significant improvement is obtained when the increasing the number of segments from 1 to 10.
+The presented value for each segment is an average over 20 optimizations values.
+Increasing the number of segments leads to an increase in the gate fidelity.
+The most significant improvement is obtained when the increasing the number of segments from 1 to 10.
 In Fig.
-5 we plot the expectation value of the fidelity for different error magnitudes.
-Empirically, we find that the fidelity for Adiabatic and Fresnel designs are very similar to the baseline (see Fig.
-9 and Appendix C for the list of the designs).
+4 we show the expectation value of the fidelity for different error magnitudes.
+Empirically, we find that the fidelity for Adiabatic and Fresnel designs are very similar to the baseline.
 We numerically observe large $N$ scaling laws for the expectation value and variance of the fidelity as functions of the number of segments:
@@
 for some constants $a$ and $b$, which we expect to be non-universal and depend on the geometry of the device, the material and the wavelength. In our example, they are $a=0.1$, and $b=0.38$.
-Figure 4: Mean Fidelity and standard deviation (shaded area) versus the number of segments for Pauli X, gate and Litium-Niobit material, assuming Normally distributed width noise with $\sigma=30$nm ($\sigma_{\%}\approx 5\%$).
+Figure 3: Mean Fidelity and standard deviation (shaded area) versus the number of segments for $\text{LiNbO}_{3}$ X gate, assuming Normally distributed width noise with $\sigma=30_{nm}$ ($\sigma_{\%}\approx 5\%$).
 We see that the mean fidelity increases and variance decreases as the number of segments grows.
 This plot was created using an ensemble average over 20 optimizations for each $N$.
-Figure 5: The expectation value of the fidelity for different number of segments, for increasing noise width.
+Figure 4: The expectation value of the fidelity for different number of segments, for increasing noise width, in $\text{LiNbO}_{3}$ X gate.
 The design is optimized for $\sigma_{\%}\approx 5\%$.
@@
 IV.2 Power Loss
-To evaluate the power loss of the design, we use Lumerical simulator to measure the transmitted light for each transition between two neighbouring segments with different widths.
-In appendix B we present a comparison between different power loss calculation methods (Fresnel coefficients, FDTD, varFDTD and EME simulations), while here we focus on the comparison between the Fresnel coefficient (up to a global factor, see appendix B for details), and the FDTD simulation result (which is considered the most accurate among the simulation methods).
+To evaluate the power loss of the design, we use Lumerical simulator to measure the transmitted light for each transition between two neighboring segments with different widths.
+In appendix B we present a comparison between different power loss calculation methods (Fresnel coefficients, FDTD, varFDTD and EME simulations), while here we focus on the comparison between the Fresnel coefficient (up to a global factor, see appendix B for details), and the FDTD simulation result, which is considered the most accurate among the simulation methods.
 Fig.
-6 shows the power loss for different number of segments, for the Hadamard gate made of Litium-Niobit material.
-While there are some differences between the Fresnel coefficients and the FDTD, their overall behaviour is very similar.
+5 shows the power loss for different number of segments, for the Hadamard gate made of $\text{LiNbO}_{3}$, each data point is averaged over 10 optimizations..
+While there are some differences between the Fresnel coefficients and the FDTD, their overall behavior is very similar.
 We find empirically that optimizing over the Fresnel coefficients leads to a reduction in the total power loss estimated by the FDTD.
 To calculate the total power loss per design, we converted the loss per transition (between segments) using Eq.
-8.
-Fig.
-6 shows the averaged result over 10 optimizations.
-Figure 6: Power loss versus number of segments for Hadamard, Litium-Niobit, using Fresnel coefficients (up to a global scale factor) in dashed line and FDTD simulation in solid line.
+19.
+Figure 5: Power loss versus number of segments for Hadamard, in $\text{LiNbO}_{3}$ based Hadamard-Gate, using Fresnel coefficients (up to a global scale factor) in dashed line and FDTD simulation in solid line.
 We see an increase in the power loss as the number of segments increases, with an improvement when employing the Fresnel and the Adiabatic smoothing methods.
@@
 6 the power density transmission cross-section curve using FDTD Lumerical simulator.
-Figure 7: Example of power density transmission cross-section curve using FDTD Lumerical simulator.
-The designs are of Silicon waveguides with $N=3$: Top - Baseline, Bottom - Fresnel-based.
-Green - Injected light in the beginning of the waveguide, blue - transmitted light to the end of the waveguide.
-The horizontal axis is the spatial cross-section of the waveguide, the vertical axis is the electromagnetic wave power density profile.
+Figure 6: An illustration comparing baseline (top) and Fresnel-based (bottom) optimizations.
+On the left, the power density transmission cross-section is depicted using the FDTD Lumerical simulator, while on the right, the geometrical configuration along the propagation axis is shown.
+The green line corresponds to the intensity profile at the waveguide’s entry, whereas the blue line indicates the final intensity profile.
+This configuration consists of 100 segments, using Silicon and a $\sqrt{X}$ gate, with a width noise standard deviation of $\sigma=30_{nm}$ ($\sigma_{\%}\approx 5\%$).
+Both designs achieve a fidelity of $0.91$.
 IV.3 Fidelity and Error Correlation
 When the correlation between the errors at different segments decreases, the composite pulse scheme for error mitigation becomes less effective YaronIdoNoiseThreshold .
-Here we analyze the impact of having a large number of segments when the error correlation decreases, where we model the error covariance matrix as (19).
+Here we analyze the impact of having a large number of segments when the error correlation decreases, where we model the error covariance matrix as Eq.
+31.
 We plot in Fig.
-8 the mean value and the standard deviation of the fidelity versus the error correlation strength $\mu$ (left, full correlation, and right, no correlation), for different numbers of segments for Pauli X gate, Silicon material.
-As expected, designs with a large number of segments show a better performance when the error correlations are small.
-We observe the same behaviour for other quantum gates.
+7 the mean value and the standard deviation of the fidelity versus the error correlation strength $\mu$ (left, full correlation, and right, no correlation), for different numbers of segments for the Pauli X gate, Silicon material.
+As expected, designs with a large number of segments show better performance when the error correlations are small.
+We observe the same behavior for other quantum gates.
 Figure 7: Mean value and standard deviation of the fidelity versus the error correlation strength (left - full correlation and right - low correlation), for different number of segments, for gate X, Silicon.
@@
 This plot was created using an ensemble average over 10 optimizations for each $N$ and optimized for $\sigma_{\%}\approx 5\%$.
-IV.4 Fidelity Versus Power Loss - An Effective Error Mitigation Region
+IV.4 Fidelity Versus Power Loss, and An Effective Error Mitigation Region
 Taking into account both the fidelity of the quantum gate and the power loss of the design, we plot in Fig.
-9 the mean fidelity versus the power loss and identify the effective error mitigation region.
-The designs are listed in table 1 in appendix C
+8 the mean fidelity versus the power loss.
+As seen, fidelity improves with a higher number of segments.
+While power loss also rises with more segments, it can be mitigated by incorporating the proposed smooth-design regularization.
+The dashed line outlines a region we term as the Effective Error Mitigation Region, as it encompasses exclusively designs that achieve both high fidelity and minimal power loss.
+All the designs presented in this figure are listed in table 1 in appendix C.
 Figure 8: Fidelity versus power loss (As calculated with the FDTD simulation).
@@
 Table 1 in appendix C gives the details of the designs.
-V Discussion and Outlook
+V Summary and Outlook
 We studied the performance of the segmented composite pulse error mitigation scheme applied to the design of photonic quantum gates, when increasing the number of segments $N$ and varying the error correlation strength $\mu$.
-We considered the fidelity and the power loss of single-qubit quantum gates realized by optical waveguides in the dual rail framework.
 We quantified the impact of increasing $N$ on the fidelity and the power loss of the quantum gates and observed certain asymptotic scaling laws.
 Empirically we found that $O(10)$ number of segments yields effective designs.
-To reduce the power loss when increasing $N$, we applied two optimization regularizations by imposing constraints on the Fresnel reflection coefficient and an adiabatic condition on the waveguides width profile.
-We have not identified a significant difference in the gate performance between the two regularizations.
-However, from practical considerations it is simpler to apply the adiabatic one since it does not require a pre-calculation of the effective refraction index in the waveguide.
-We studied the effect of varying the correlation strength between the errors of different segments.
-Large correlations between the errors imply better gate fidelity, since the different segments can effectively compensate the accumulated errors.
-We showed that a large number of segments allows to overcome the problem of weak error correlations.
-There are several directions to follow.
-One is to analyze the continuum limit of the large $N$ designs that result in smooth designs.
-A second one is to validate our error model against experimental data.
-A third direction to follow is to study the robust designs for QIP systems that include several or many quantum gates.
+To minimize power loss with increasing $N$, we applied two optimization regularization: one imposing constraints on the Fresnel reflection coefficient and the other introducing an adiabatic condition on the waveguide width profile.
+Both approaches yielded comparable gate performance, although the adiabatic condition proves to be slightly more convenient within the optimization context.
+We also examined the impact of varying correlation strength between errors in different segments.
+Our results showed that stronger correlations improve gate fidelity, as the segments can better compensate for accumulated errors.
+Notably, a large number of segments mitigates the challenges associated with weak error correlations.
+There are several natural research directions to follow.
+First, while our results are specifically tailored to dual-rail photonic gates, the insights gained from this study are expected to generalize across various noise models and error frameworks, including those with different error correlations.
+These findings hold potential for application to other quantum systems and platforms, providing valuable guidance for error mitigation strategies in broader quantum contexts.
+Another important direction to follow is to study robust designs for QIP systems that include several or many quantum gates.
+In such cases, the optimization framework is likely to find robust designs that are better than a simple combination of individually robust single-qubit gates.
 Acknowledgements We would like to thank J.
@@
 Piasetzky for their assistance in extracting parameters using Lumerical simulations.
-This work is supported in part by the Israeli Science Foundation Excellence Center, the US-Israel Binational Science Foundation, and the Israel Ministry of Science.
-Appendix A Single Qubit Gates
-The time evolution of a two-level systems (qubit) is governed by the Schrödinger equation:
-$i\partial_{t}\left(\begin{array}[]{l}c_{1}(t)\\ c_{2}(t)\end{array}\right)=\left(\begin{array}[]{cc}-\Delta(t)&\kappa^{*}(t)\\ \kappa(t)&\Delta(t)\end{array}\right)\left(\begin{array}[]{l}c_{1}(t)\\ c_{2}(t)\end{array}\right)\ ,$ (28)
-where $c_{1}(t)$ and $c_{2}(t)$ are the probability amplitudes at time $t$ of being in the states $|0\rangle$ and $|1\rangle$, respectively, $\kappa(t)$ is the (complex) Rabi frequency, $\Delta(t)$ is the (real) mode-mismatch, and we set $\hbar=1$.
-The unitary propagator of such a system is:
-$U(t,0)=\mathcal{T}\left[\exp\left[-i\int_{0}^{t}\left(\begin{array}[]{cc}-\Delta\left(t^{\prime}\right)&\kappa^{*}\left(t^{\prime}\right)\\ \kappa\left(t^{\prime}\right)&\Delta\left(t^{\prime}\right)\end{array}\right)\right]dt^{\prime}\right]\ .$ (31)
-When $\kappa$ and $\Delta$ are time independent, the propagator simplifies to
-$U(t,0)=\exp\left[-it\left(\begin{array}[]{cc}-\Delta&\kappa^{*}\\ \kappa&\Delta\end{array}\right)\right]\ .$ (34)
-Physical systems that exhibit this $SU(2)$ dynamics implement one-qubit gates as well as probabilistic two-qubit gates.
-In silicon photonics, the analog of the time is the propagation axis, and the coupling coefficient $\kappa$ is real.
-Therefore, for a single segment of length $z$:
-$U(\kappa,\Delta,z)=\exp\left[-iz\left(\kappa X-\Delta Z\right)\right]\ ,$ (35)
-where $X$ and $Z$ are the Pauli matrices. We can rewrite the $SU(2)$ quantum gate as:
-$U(\kappa,\Delta,z)=\exp\left[-iz\Omega_{g}\left(\frac{\kappa}{\Omega_{g}},0,\frac{\Delta}{\Omega_{g}}\right)\cdot\vec{\sigma}\right]$ (36)
-$=\cos\left(z\Omega_{g}\right)I-\frac{\kappa}{\Omega_{g}}\sin\left(z\Omega_{g}\right)X-\frac{\Delta}{\Omega_{g}}\sin\left(z\Omega_{g}\right)Z\ ,$ (37)
-where $\Omega_{g}\equiv\sqrt{\kappa^{2}+\Delta^{2}}$.
+This work is supported in part by the Israeli Science Foundation Excellence Center No.
+2312/21, the US-Israel Binational Science Foundation, and the Israel Ministry of Science.
+Appendix A Bloch Sphere Trajectories
+Fig. 9 shows a trajectory comparison for two different cases.
+Figure 9: Bloch Sphere diagram of two configurations in Silicon.
+Left - Hadamard gate with 10 segments, initial state is $|0\rangle$ and final state is |+ket|$\rangle$| + ⟩.
+Right - Pauli X gate with 50 segments, initial state is $|+\rangle_{y}$ and final state is $|-\rangle_{y}$.
 Appendix B Power Loss Simulation Comparison
@@
 We consider the FDTD the most accurate simulation method and this is the reason we present this method along with the Fresnel coefficients.
-The Fresnel coefficient approximation of the power loss compared to the simulation, are related by a global factor of about 5 (see 6 in the main text).
+The Fresnel coefficient approximation of the power loss compared to the simulation, are related by a global factor of about 5 (see Fig.
+5 in the main text).
 Figure 10: Our power loss simulation using EME (top-left), Fresnel coefficients (top-right), varFDTD (bottom-left) and FDTD (bottom-right).
-The simulation evaluated the power loss due to a single abrupt transition from width $W_{1}$ (x axis in nm) to width $W_{2}$ (y axis in nm), for Silicon and $H=220nm$.
+The simulation evaluated the power loss due to a single abrupt transition from width $W_{0}$ (x axis in nm) to width $W_{1}$ (y axis in nm), for Silicon and $H=220nm$.
 Appendix C Table of designs